An astronaut on the moon tosses a ball upward; by how much will the height of the ball exceed 100 feet if the ball's height follows: h(t)=-2.6t^2+32.1t+5.5
Answer by stanbon(57387)

An astronaut on the moon tosses a ball upward; by how much will the height of the ball exceed 100 feet if the ball's height follows: h(t)=-2.6t^2+32.1t+5.5
-----------------
Maximum height occurs when t = -b/2a = -32.1/(2*-2.6) = 6.173 seconds
---------------
h(6.173) = 104.58 ft.
-----------------------
The ball will exceed 100 ft. in height by 4.58 feet.
============
Cheers,
Stan H.